# Weighted sampling recovery of functions with mixed smoothness

### v3

@@ Abstract
We studied linear weighted sampling algorithms and their optimality for approximate recovery of functions with mixed smoothness on $\mathbb{R}^{d}$ from a set of $n$ their sampled values. Functions to be recovered are in weighted Sobolev spaces $W^{r}_{p,w}(\mathbb{R}^{d})$ of mixed smoothness, and the approximation error is measured by the norm of the weighted Lebesgue space $L_{q,w}(\mathbb{R}^{d})$. Here, the weight $w$ is a tensor-product Freud-type weight. The optimality of linear sampling algorithms is investigated in terms of sampling $n$-widths. We constructed linear sampling algorithms on sparse grids of sampled points which form a step hyperbolic cross in the function domain, and which give upper bounds for the corresponding sampling $n$-widths. We proved that in the one-dimensional case, these algorithms realize the right convergence rate of the $n$-sampling widths.
@@ 1 Introduction
Thus, the spaces $L_{p}(\Omega;\gamma)$ and $W^{r}_{p}(\Omega;\gamma)$ coincide with $L_{p,w}(\Omega)$ and $W^{r}_{p,w}(\Omega)$, where $w:=\left(w_{\operatorname{g}}\right)^{1/p}$ for a fixed $1\leq p<\infty$. For a normed space $X$ of functions on $\Omega$, the boldface $\boldsymbol{X}$} denotes the unit ball in $X$.
@@
There is a large number of works devoted to the problem of (unweighted) linear sampling recovery of functions having a mixed smoothness on a compact domain, see for survey and bibliography in [7], [15], [17], [3]. Smolyak sparse-grid sampling algorithms for compact domains (cube or torus) have been widely and efficiently used in both approximation theory and numerical analysis, especially, in sampling recovery for functions having a mixed smoothness. A huge body of papers investigated various aspects on approximation and numerical implementation by Smolyak sparse-grid sampling algorithms. The reader can consult [1], [7], [15], [17] for a survey and bibliography.
The optimal convergence rate of sampling $n$-widths is one of central problems in sampling recovery of functions having a mixed smoothness. We refer the reader to [3, 7] for a survey and bibliography on the optimal convergence rate of the sampling $n$-widths $\varrho_{n}\big{(}{\boldsymbol{W}}^{r}_{p}({\mathbb{T}}^{d}),L_{q}({\mathbb{T}}^{d})\big{)}$, where ${\mathbb{T}}^{d}$ denotes the $d$-dimensional torus. It is interesting to notice that the right convergence of these widths in the cases $1<p<q\leq 2$, $2\leq p<q<\infty$ and $p=2,\ q=\infty$ can be achieved by Smolyak sparse-grid sampling algorithms which are asymptotically optimal, and besides of that so far we do not know any other type of asymptotically optimal algorithm. From results of [9] on inequality between sampling and Kolmogorov $n$-widths and known results on relevant Kolmogorov $n$-widths, one can deduce in a non-constructive way the right convergence rate of $\varrho_{n}({\boldsymbol{W}}^{r}_{p}({\mathbb{T}}^{d}),L_{q}({\mathbb{T}}^{d}))$ for the case $1<q\leq 2\leq p\leq\infty$ (see [3]).
The problem of optimal sampling recovery of functions on ${\mathbb{R}}^{d}$ equipped with Gaussian measure has been investigated in [6]. We proved in a non-constructive way the right convergence rate of the sampling $n$-widths
@@
which is obtained by using inequalities between sampling widths and Kolmogorov widths and the right convergence rate of $d_{n}\big{(}{\boldsymbol{W}}^{r}_{p}({\mathbb{R}}^{d};\gamma),L_{q}({\mathbb{R}}^{d};\gamma)\big{)}$ proven in the same paper [6].
In the present paper, we are interested in constructing sparse-grid linear sampling algorithms for approximate recovery of functions with mixed smoothness on $\mathbb{R}^{d}$ in the following setting: functions to be recovered are in weighted Sobolev spaces $W^{r}_{p,w}(\mathbb{R}^{d})$ of mixed smoothness $r$, and the approximation error is measured by the norm of the weighted Lebesgue space $L_{q,w}(\mathbb{R}^{d})$. The parameters $p$ and $q$ may be different and range in $1,\infty]$. The asymptotic optimality of linear sampling algorithms is investigated in terms of the relevant sampling widths. We construct sparse-grid sampling algorithms which give upper bounds of the sampling $n$-widths $\varrho_{n}(\boldsymbol{W}^{r}_{p,w}(\mathbb{R}^{d}),L_{q,w}({\mathbb{R}}^{d}))$. In, the case $d=1$, these algorithms will realize the right convergence rate of the sampling widths. In particular, we develop a counterpart of Smolyak sparse-grid sampling algorithms which realize the upper bounds.
@@
In the one-dimensional case, we prove the right convergence rate
@@
Some results on univariate weighted interpolation from [12] are employed in proving the upper bounds (1.10). The Smolyak sampling algorithms performing these upper bounds are constructed on sparse grids of sampled points which form a step hyperbolic cross in the function domain ${\mathbb{R}}^{d}$. It is remarkable to notice that these sparse grids are completely different from classical hyperbolic crosses in frequency domains and Smolyak sparse grids in compact function domains (see Figures 1 and 2). To prove the lower bounds in (1.11) and in (1.12) in the case $p<q$ we adopt a traditional technique to construct for arbitrary $n$ sampled points a fooling function vanishing at these points. The lower bounds in (1.11) and in (1.12) in the case $p\geq q$ are derived from the lower bound of Kolmogorov widths, a Bernstein-type inequality by using a discretization technique.
@@
The paper is organized as follows. In Section 2, we prove the right convergence rate of $\varrho_{n}({\boldsymbol{W}}^{r}_{p,w}({\mathbb{R}}),L_{q,w}({\mathbb{R}}))$ and construct asymptotically optimal linear sampling algorithms. In Section 3, we prove upper and lower bounds of $\varrho_{n}({\boldsymbol{W}}^{r}_{p,w}({\mathbb{R}}^{d}),L_{q,w}({\mathbb{R}}^{d}))$ for $d\geq 2$, and construct linear sampling algorithms which give the upper bound.
Notation. Denote ${\boldsymbol{x}}=:\left(x_{1},...,x_{d}\right)$ for ${\boldsymbol{x}}\in{\mathbb{R}}^{d}$; ${\boldsymbol{1}}:=(1,...,1)\in{\mathbb{R}}^{d}$; $|{\boldsymbol{x}}|_{p}:=\left(\sum_{j=1}^{d}|x_{j}|^{p}\right)^{1/p}$ $(1\leq p<\infty)$ and $|{\boldsymbol{x}}|_{\infty}:=\max_{1\leq j\leq d}|x_{j}|$ with the abbreviation: $|{\boldsymbol{x}}|:=|{\boldsymbol{x}}|_{2}$. For ${\boldsymbol{x}},{\boldsymbol{y}}\in{\mathbb{R}}^{d}$, the inequality ${\boldsymbol{x}}\leq{\boldsymbol{y}}$ (${\boldsymbol{x}}<{\boldsymbol{y}}$) means $x_{i}\leq y_{i}$ ($x_{i}<y_{i}$) for every i=1,,di=1,...,d. For $x\in{\mathbb{R}}$, denote $\operatorname{sign}(x):=1$ if $x\geq 0$, and $\operatorname{sign}(x):=-1$ if $x<0$. We use letters $C$ and $K$ to denote general positive constants which may take different values. For the quantities $A_{n}(f,{\boldsymbol{k}})$ and $B_{n}(f,{\boldsymbol{k}})$ depending on $n\in{\mathbb{N}}$, $f\in W$, ${\boldsymbol{k}}\in{\mathbb{Z}}^{d}$, we write $A_{n}(f,{\boldsymbol{k}})\ll B_{n}(f,{\boldsymbol{k}})$, $f\in W$, ${\boldsymbol{k}}\in{\mathbb{Z}}^{d}$ ($n\in{\mathbb{N}}$ is specially dropped), if there exists some constant $C>0$ independent of $n,f,{\boldsymbol{k}}$ such that $A_{n}(f,{\boldsymbol{k}})\leq CB_{n}(f,{\boldsymbol{k}})$ for all $n\in{\mathbb{N}}$, $f\in W$, ${\boldsymbol{k}}\in{\mathbb{Z}}^{d}$ (the notation $A_{n}(f,{\boldsymbol{k}})\gg B_{n}(f,{\boldsymbol{k}})$ has the obvious opposite meaning), and $A_{n}(f,{\boldsymbol{k}})\asymp B_{n}(f,{\boldsymbol{k}})$ if $S_{n}(f,{\boldsymbol{k}})\ll B_{n}(f,{\boldsymbol{k}})$ and $B_{n}(f,{\boldsymbol{k}})\ll S_{n}(f,{\boldsymbol{k}})$. Denote by $|G|$ the cardinality of the set $G$. For a Banach space $X$, denote by the boldface $\boldsymbol{X}$} the unit ball in $X$.
@@ 2 Univariate sampling recovery
In this section, for the one-dimensional case, we prove the right convergence rate of sampling $n$-widths $\varrho_{n}\big{(}{\boldsymbol{W}}^{r}_{p,w}({\mathbb{R}}),L_{q,w}({\mathbb{R}})\big{)}$ and construct asymptotically optimal linear sampling linear algorithms for $1<p<\infty$ and $1\leq q\leq\infty$. We also give some upper and lower bounds of $\varrho_{n}\big{(}{\boldsymbol{W}}^{r}_{\infty,w}({\mathbb{R}}),L_{q,w}({\mathbb{R}})\big{)}$ for $1\leq q\leq\infty$.
@@
$\big{\|}\Delta_{k}^{I}(f)\big{\|}_{L_{q,w}({\mathbb{R}})}\leq C2^{-r_{\lambda,p,q}k}\|f\|_{W^{r}_{p,w}({\mathbb{R}})},\ \ k\in{\mathbb{N}}_{0}.$ (2.20)
Proof. Let $f\in W^{r}_{p,w}({\mathbb{R}})$. Since $\Delta_{k}^{I}f\in{\mathcal{P}}_{m_{k}+1}$ by the claims (iii) and (iv) of Lemma 2.3 we have that
$\big{\|}\Delta_{k}^{I}(f)\big{\|}_{L_{q,w}({\mathbb{R}})}\ll 2^{\delta_{\lambda,p,q}k}\big{\|}\Delta_{k}^{I}(f)\big{\|}_{L_{p,w}({\mathbb{R}})},\ \ k\in{\mathbb{N}}_{0}.$ (2.21)
@@
$\big{\|}\Delta_{k}^{I}(f)\big{\|}_{L_{p,w}({\mathbb{R}})}$ $\leq\big{\|}f-I_{2j(m_{k})}(f)\big{\|}_{L_{p,w}({\mathbb{R}})}+\big{\|}f-I_{2j(m_{k-1})}(f)\big{\|}_{L_{p,w}({\mathbb{R}})}$
@@
$\begin{split}\big{\|}f-I_{m_{k(n)}}(f)\big{\|}_{L_{q,w}({\mathbb{R}})}&=\Bigg{\|}\sum_{k>k(n)}\Delta_{k}^{I}(f)\Bigg{\|}_{L_{q,w}({\mathbb{R}})}\leq\sum_{k>k(n)}\big{\|}\Delta_{k}^{I}(f)\big{\|}_{L_{q,w}({\mathbb{R}})}\\ &\ll\sum_{k>k(n)}2^{-r_{\lambda,p,q}k}\|f\|_{W^{r}_{p,w}({\mathbb{R}}^{d})}\\ &\leq 2^{-r_{\lambda,p,q}k(n)}\sum_{k>k(n)}2^{-r_{\lambda,p,q}(k-k(n))}\ll n^{-r_{\lambda,p,q}},\end{split}$ (2.26)
@@
${\mathcal{B}}^{*}_{N(m),p}:=\left\{\varphi\in{\mathcal{P}}^{*}_{m+1}:\|\varphi\|_{L_{p,w}({\mathbb{R}})}\leq 1\right\}.$
@@
$\big{\|}\Delta_{k}^{I}(f)\big{\|}_{L_{q,w}({\mathbb{R}})}\ll 2^{-r_{\lambda,\infty,q}k}k\,\|f\|_{W^{r}_{\infty,w}({\mathbb{R}})},\ \ k\in{\mathbb{N}}_{0}.$ (2.40)
@@ 3 Sparse-grid sampling recovery in high dimension
Based on the operators $\Delta_{k}^{I},\ k\in{\mathbb{N}}_{0}$, defined in (2.18), we construct sampling operators on ${\mathbb{R}}^{d}$ by using the well-known Smolyak algorithm. For convenience, with an abuse we make use of the notation:
@@
Figure 1: Different hyperbolic crosses (HC)($d=2$)
Figure 2: A classical Smolyak grid ($d=2$)

### v4

@@ Abstract
We studied linear weighted sampling algorithms and their optimality for approximate recovery of functions with mixed smoothness on $\mathbb{R}^{d}$ from a set of $n$ their sampled values. Functions to be recovered are in weighted Sobolev spaces $W^{r}_{p,w}(\mathbb{R}^{d})$ of mixed smoothness, and the approximation error is measured by the norm of the weighted Lebesgue space $L_{q,w}(\mathbb{R}^{d})$. Here, the weight $w$ is a tensor-product Freud-type weight. The optimality of linear sampling algorithms is investigated in terms of sampling $n$-widths. We constructed linear sampling algorithms on sparse grids of sampled points which form a step hyperbolic cross in the function domain, and which give upper bounds for the corresponding sampling $n$-widths. We proved that in the one-dimensional case, these algorithms realize the exact convergence rate of the $n$-sampling widths.
@@ 1 Introduction
Thus, the spaces $L_{p}(\Omega;\gamma)$ and $W^{r}_{p}(\Omega;\gamma)$ coincide with $L_{p,w}(\Omega)$ and $W^{r}_{p,w}(\Omega)$, where $w:=\left(w_{\operatorname{g}}\right)^{1/p}$ for a fixed $1\leq p<\infty$. Throughout this paper, we use ${\boldsymbol{W}}^{r}_{p,w}(\Omega)$ and ${\boldsymbol{W}}^{r}_{p}(\Omega;\gamma)$ to denote the unit ball in the spaces $W^{r}_{p,w}(\Omega)$ and $W^{r}_{p}(\Omega;\gamma)$, respectively.
@@
There is a large number of works devoted to the problem of (unweighted) linear sampling recovery of functions having a mixed smoothness on a compact domain. Smolyak sparse-grid sampling algorithms for compact domains (cube or torus) have been widely and efficiently used in both approximation theory and numerical analysis, especially, in sampling recovery for functions having a mixed smoothness. The reader can consult [1, 3, 7, 15, 17] for a survey and bibliography. The optimal convergence rate of sampling $n$-widths is one of central problems in sampling recovery of functions having a mixed smoothness. We refer the reader to [3, 7] for a survey and bibliography on the optimal convergence rate of the sampling $n$-widths $\varrho_{n}\big({\boldsymbol{W}}^{r}_{p}({\mathbb{T}}^{d}),L_{q}({\mathbb{T}}^{d})\big)$, where ${\boldsymbol{W}}^{r}_{p}({\mathbb{T}}^{d})$ is the unit ball in the Sobolev space $W^{r}_{p}({\mathbb{T}}^{d})$ of functions on the $d$-dimensional torus ${\mathbb{T}}^{d}$ having mixed smoothness $r$. It is noteworthy that the recent results of [9] on inequality between the sampling and Kolmogorov $n$-widths of the unit ball in a reproducing kernel Hilbert subspace immediately yield in a nonconstructive way the exact convergence rate of $\varrho_{n}({\boldsymbol{W}}^{r}_{2}({\mathbb{T}}^{d}),L_{2}({\mathbb{T}}^{d}))$. This solved the outstanding Open Problem 4.1 in [7] for the particular case $p=2$.
The problem of optimal sampling recovery of functions on ${\mathbb{R}}^{d}$ equipped with Gaussian measure has been investigated in [6]. We proved in a non-constructive way the exact convergence rate of the sampling $n$-widths
@@
which is obtained by using the inequalities between sampling widths and Kolmogorov widths [9] mentioned above, and the exact convergence rate of $d_{n}\big({\boldsymbol{W}}^{r}_{p}({\mathbb{R}}^{d};\gamma),L_{q}({\mathbb{R}}^{d};\gamma)\big)$ proven in the same paper [6]. A related problem on hyperbolic cross weighted polynomial approximation based on based on de la Vallée Poussin sums of the relevant orthonormal polynomial expansion, has been treated in [4].
In the present paper, we are interested in constructing sparse-grid linear sampling algorithms for approximate recovery of functions with mixed smoothness on $\mathbb{R}^{d}$ in the following setting: functions to be recovered are in weighted Sobolev spaces $W^{r}_{p,w}(\mathbb{R}^{d})$ of mixed smoothness $r$, and the approximation error is measured by the norm of the weighted Lebesgue space $L_{q,w}(\mathbb{R}^{d})$. The parameters $p$ and $q$ may be different and range in $1,\infty]$. The asymptotic optimality of linear sampling algorithms is investigated in terms of the relevant sampling widths. We construct sparse-grid sampling algorithms which give upper bounds of the sampling $n$-widths $\varrho_{n}(\boldsymbol{W}^{r}_{p,w}(\mathbb{R}^{d}),L_{q,w}({\mathbb{R}}^{d}))$. In, the case $d=1$, these algorithms will realize the exact convergence rate of the sampling widths. In particular, we develop a counterpart of Smolyak sparse-grid sampling algorithms which realize the upper bounds.
@@
In the one-dimensional case, we prove the exact convergence rate
@@
Some results on univariate weighted interpolation from [12] are employed in proving the upper bounds (1.10). The Smolyak sampling algorithms performing these upper bounds are constructed on sparse grids of sampled points which form a step hyperbolic cross in the function domain ${\mathbb{R}}^{d}$. This development of the construction of sparse grids and the associated algorithms was proposed first by the author [2] for numerical weighted integration of functions having weighted mixed smoothness. It is remarkable to notice that these sparse grids are completely different from classical hyperbolic crosses in frequency domains and Smolyak sparse grids in compact function domains (see Figures 1 and 2). To prove the lower bounds in (1.11) and (1.12) in the case $p<q$ we adopt a traditional technique to construct for arbitrary $n$ sampled points a fooling function vanishing at these points. The lower bounds in (1.11) and (1.12) the case $p\geq q$ are derived from the lower bound of Kolmogorov widths, a Bernstein-type inequality by using a discretization technique.
@@
The paper is organized as follows. In Section 2, we prove the exact convergence rate of $\varrho_{n}({\boldsymbol{W}}^{r}_{p,w}({\mathbb{R}}),L_{q,w}({\mathbb{R}}))$ and construct asymptotically optimal linear sampling algorithms. In Section 3, we prove upper and lower bounds of $\varrho_{n}({\boldsymbol{W}}^{r}_{p,w}({\mathbb{R}}^{d}),L_{q,w}({\mathbb{R}}^{d}))$ for $d\geq 2$, and construct linear sampling algorithms which give the upper bound.
Notation. Denote ${\boldsymbol{x}}=:\left(x_{1},...,x_{d}\right)$ for ${\boldsymbol{x}}\in{\mathbb{R}}^{d}$; ${\boldsymbol{1}}:=(1,...,1)\in{\mathbb{R}}^{d}$; $|{\boldsymbol{x}}|_{p}:=\left(\sum_{j=1}^{d}|x_{j}|^{p}\right)^{1/p}$ $(1\leq p<\infty)$ and $|{\boldsymbol{x}}|_{\infty}:=\max_{1\leq j\leq d}|x_{j}|$ with the abbreviation: $|{\boldsymbol{x}}|:=|{\boldsymbol{x}}|_{2}$. For ${\boldsymbol{x}},{\boldsymbol{y}}\in{\mathbb{R}}^{d}$, the inequality ${\boldsymbol{x}}\leq{\boldsymbol{y}}$ (${\boldsymbol{x}}<{\boldsymbol{y}}$) means $x_{i}\leq y_{i}$ ($x_{i}<y_{i}$) for every i=1,,di=1,...,d. For $x\in{\mathbb{R}}$, denote $\operatorname{sign}(x):=1$ if $x\geq 0$, and $\operatorname{sign}(x):=-1$ if $x<0$. We use letters $C$ and $K$ to denote general positive constants which may take different values. For the quantities $A_{n}(f,{\boldsymbol{k}})$ and $B_{n}(f,{\boldsymbol{k}})$ depending on $n\in{\mathbb{N}}$, $f\in W$, ${\boldsymbol{k}}\in{\mathbb{Z}}^{d}$, we write $A_{n}(f,{\boldsymbol{k}})\ll B_{n}(f,{\boldsymbol{k}})$, $f\in W$, ${\boldsymbol{k}}\in{\mathbb{Z}}^{d}$ ($n\in{\mathbb{N}}$ is specially dropped), if there exists some constant $C>0$ independent of $n,f,{\boldsymbol{k}}$ such that $A_{n}(f,{\boldsymbol{k}})\leq CB_{n}(f,{\boldsymbol{k}})$ for all $n\in{\mathbb{N}}$, $f\in W$, ${\boldsymbol{k}}\in{\mathbb{Z}}^{d}$ (the notation $A_{n}(f,{\boldsymbol{k}})\gg B_{n}(f,{\boldsymbol{k}})$ has the obvious opposite meaning), and $A_{n}(f,{\boldsymbol{k}})\asymp B_{n}(f,{\boldsymbol{k}})$ if $S_{n}(f,{\boldsymbol{k}})\ll B_{n}(f,{\boldsymbol{k}})$ and $B_{n}(f,{\boldsymbol{k}})\ll S_{n}(f,{\boldsymbol{k}})$. Denote by $|G|$ the cardinality of the set $G$.
@@ 2 Univariate sampling recovery
In this section, for the one-dimensional case, we prove the exact convergence rate of sampling $n$-widths $\varrho_{n}\big({\boldsymbol{W}}^{r}_{p,w}({\mathbb{R}}),L_{q,w}({\mathbb{R}})\big)$ and construct asymptotically optimal linear sampling linear algorithms for $1<p<\infty$ and $1\leq q\leq\infty$. We also give some upper and lower bounds of $\varrho_{n}\big({\boldsymbol{W}}^{r}_{\infty,w}({\mathbb{R}}),L_{q,w}({\mathbb{R}})\big)$ for $1\leq q\leq\infty$.
@@
$\big\|\Delta_{k}(f)\big\|_{L_{q,w}({\mathbb{R}})}\leq C2^{-r_{\lambda,p,q}k}\|f\|_{W^{r}_{p,w}({\mathbb{R}})},\ \ k\in{\mathbb{N}}_{0}.$ (2.20)
Proof. Let $f\in W^{r}_{p,w}({\mathbb{R}})$. Since $\Delta_{k}f\in{\mathcal{P}}_{m_{k}+1}$ by the claims (iii) and (iv) of Lemma 2.3 we have that
$\big\|\Delta_{k}(f)\big\|_{L_{q,w}({\mathbb{R}})}\ll 2^{\delta_{\lambda,p,q}k}\big\|\Delta_{k}(f)\big\|_{L_{p,w}({\mathbb{R}})},\ \ k\in{\mathbb{N}}_{0}.$ (2.21)
@@
$\big\|\Delta_{k}(f)\big\|_{L_{p,w}({\mathbb{R}})}$ $\leq\big\|f-I_{2j(m_{k})}(f)\big\|_{L_{p,w}({\mathbb{R}})}+\big\|f-I_{2j(m_{k-1})}(f)\big\|_{L_{p,w}({\mathbb{R}})}$
@@
$\begin{split}\big\|f-I_{m_{k(n)}}(f)\big\|_{L_{q,w}({\mathbb{R}})}&=\Bigg\|\sum_{k>k(n)}\Delta_{k}(f)\Bigg\|_{L_{q,w}({\mathbb{R}})}\leq\sum_{k>k(n)}\big\|\Delta_{k}(f)\big\|_{L_{q,w}({\mathbb{R}})}\\ &\ll\sum_{k>k(n)}2^{-r_{\lambda,p,q}k}\|f\|_{W^{r}_{p,w}({\mathbb{R}}^{d})}\\ &\leq 2^{-r_{\lambda,p,q}k(n)}\sum_{k>k(n)}2^{-r_{\lambda,p,q}(k-k(n))}\ll n^{-r_{\lambda,p,q}},\end{split}$ (2.26)
@@
${\mathcal{B}}^{*}_{N(m),p}:=\left\{\varphi\in{\mathcal{P}}^{*}_{m+1}:\|\varphi\|_{L_{p,w}({\mathbb{R}})}\leq 1\right\}$
@@
$\big\|\Delta_{k}(f)\big\|_{L_{q,w}({\mathbb{R}})}\ll 2^{-r_{\lambda,\infty,q}k}k\,\|f\|_{W^{r}_{\infty,w}({\mathbb{R}})},\ \ k\in{\mathbb{N}}_{0}.$ (2.40)
@@ 3 Sparse-grid sampling recovery in high dimension
Based on the operators $\Delta_{k},\ k\in{\mathbb{N}}_{0}$, defined in (2.18), we construct sampling operators on ${\mathbb{R}}^{d}$ by using the well-known Smolyak algorithm. For convenience, with an abuse we make use of the notation:
@@
Figure 1: Different hyperbolic crosses (HC)($d=2$) borrowed from [2]
Figure 2: A classical Smolyak grid ($d=2$) borrowed from [2]